HOT WATER EMISSION IN THE NEAR INFRARED

\underline{K. TERESZCHUK}, T. HIRAO, and P. F. BERNATH, Department of Chemistry, University of Waterloo, Waterloo, Ontario, CANADA N2L 3G1.

Hot water lines are found in a variety of sources including hydrocarbon flames and astronomical objects such as brown dwarfs. Thus far, our laboratory has reported on high resolution emission spectra of water vapor at 1550 ^{∘}C in the region of 400 − 6000 cm^{−1}. In our current investigation, we have now recorded our laboratory hot water spectra from 5000 − 9000 cm^{−1} with a Bruker IFS 120 HR spectrometer. The spectra will be compared with absorption spectra of water in sunspots in the same spectral region.